OpenStudy (madhatter):

can u Solve the equation for y. 5x – 4y = 12

OpenStudy (anonymous):

5x-4y=12 -5x+5x-4y=12-5x -4y=12-5x divide by -4 on both sides u get y= 3-5x

OpenStudy (madhatter):

thanks

OpenStudy (anonymous):

ur welcome

OpenStudy (ja1):

Grrrrr, I'm late <_<

OpenStudy (madhatter):

no ur not the frist one didnt make sence

OpenStudy (ja1):

haha i see well what do you not understand?

OpenStudy (madhatter):

well the equation was not right it just didnt look like the right answer

OpenStudy (ja1):

Well funy thing is I studied this like 3 months ago but I don't remember xD but it actually does seem correct to be honest it has the rght format.

OpenStudy (madhatter):

oh i c its ment for graphing thats y i wasnt so sure

OpenStudy (ja1):

OHHHHH

OpenStudy (ja1):

ok that changes it all

OpenStudy (ja1):

not all but some of it :D

OpenStudy (madhatter):

ya lol

OpenStudy (madhatter):

i am doing connections academy so i dont get much help with understanding stuff

OpenStudy (ja1):

sorry my wifi just went crazy lol

OpenStudy (ja1):

so you are finding coordinates right?

OpenStudy (madhatter):

its fine

OpenStudy (ja1):

ok so if you are finding coordinates then we this got alot simpler one sec.

OpenStudy (madhatter):

kk

OpenStudy (ja1):

I think I may have mental problems <_< here I am trying to solve what I belive is a system of equations (>_<)

OpenStudy (madhatter):

lol haha i have that prob alot the mental prob that is

OpenStudy (ja1):

Okay so here we go I may take a while I must have it all neet, OCD and stuff xD

OpenStudy (madhatter):

ya lol haha

OpenStudy (ja1):

5x – 4y = 12 5(0) - 4y = 12

OpenStudy (ja1):

to solve for Y we must substitue x iwith 0 so 5 is taken out

OpenStudy (madhatter):

ok

OpenStudy (ja1):

so we now have -4y = 12 since the negative sign stays with 4

OpenStudy (ja1):

now solve that by isolating the variable

OpenStudy (madhatter):

ok

OpenStudy (madhatter):

thanks makes sence

OpenStudy (ja1):

yupp no problem

OpenStudy (ja1):

viceversa for X

OpenStudy (madhatter):

ok that helps alot

OpenStudy (ja1):

I hope so :) Now I must go many more questions to solve :D

OpenStudy (madhatter):

ya lol i was trying to help with an prob be she didnt get it

OpenStudy (ja1):

HaHa yah sometimes it's best to just leave xD